A rectangular Persian carpet has a perimeter of 196 in. The length of the carpet is 18in more than the width. What are the dimensions of the carpet?
Answer by JulietG(1812)

Perimeter = 2length + 2width
P = 2L + 2W
P = 196
L = W + 18
Substitute the known value of L into the equation
196 = 2(W+18) + 2W
196 = 2W + 36 + 2W
196 = 4W + 36
Subtract 36 from each side
160 = 4W
Divide each side by 4
40 = W
If the width is 40, then the Length is 40+18, or 58
Check:
196 = 2(58) + 2(40)
196 = 116 + 80
196 = 196
Success!